GAS PHASE ELECTRONIC PHOTODISSOCIATION SPECTRA OF COPPER NITRATE CLUSTER IONS

SYDNEY H. KAUFMAN, \underline{CASEY R. CHRISTOPHER} and J. MATHIAS WEBER, JILA, NIST and Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309.

We report UV photodissociation action spectra of copper nitrate cluster anions Cu(NO_{3})_{3}^{−}. The parent ion decays upon electronic excitation by loss of neutral NO_{3} fragment. THe gas phase spectrum is similar to but slightly more structured than UV absorption spectra of solutions of the parent salt Cu(NO_{3})_{2}. We discuss our experimental results in the framework of quantum chemical calculations.